OpenStudy (mortonsalt):

Evaluate using polar coordinates: f(x,y)=4, and the region R is given by r=sin(2theta) in the first quadrant.

OpenStudy (hacer):

according to first quadrant we have \[0\le \theta \le \Pi/2\] then for the integral \[\int\limits_{0}^{\Pi/2}\int\limits_{0}^{\sin(2\theta)}4rdrd \theta \] now taking 4 out of the integral firs I'll solve according to r's integral such that \[\int\limits_{0}^{\sin(2\theta)}rdr=r^2/2 \] then putting the equation in order to upper and lower bound .I have \[\sin^2(2\theta)/2-0=\sin^2(2\theta)/2\] I know \[\sin^2(2\theta)/2=1-\cos(4\theta)/2\times1/2=1-\cos(4\theta)/4\] Then I have new integral such that \[4\int\limits_{0}^{\Pi/2}(1-\cos(4\theta))/4 d \theta \] taking 1/4 out of the integral I get \[\int\limits_{0}^{\Pi/2}(1-\cos(4\theta))d \theta \] \[=\theta-\sin(4\theta)/4\] putting upper and lower bound in order to \[\theta \] the I get \[[\Pi /2-\sin(4\times \Pi/2)/4]-[0-\sin(0)/4]=\Pi/2\] so the result is \[\Pi/2\]